# Contextual modulation of language comprehension in a dynamic neural model of lexical meaning

### v1

@@ 2.2 Model simulations
Existing empirical results have also demonstrated that individual variation in the magnitude of the by-context difference in acceptability is predicted by AQ scores, such that individuals with higher AQ scores show a reduced influence of context. We model individual variation in AQ by varying the parameter , which controls the magnitude of field coupling. Stronger field coupling is consistent with a greater degree of system-level expectations. In other words, given some interpretation on one semantic dimension (e.g., control asymmetry (CA)), individuals can vary in the degree to which they expect a corresponding interpretation on a related semantic dimension (e.g., connectedness (conn)). We posit that individuals with higher AQ scores are more influenced by system-level expectations, i.e. stronger field coupling. With stronger system-level expectations, the processing system is more rigid, and less influenced by real-time signals. This is consistent with existing findings relating AQ scores to linguistic behavior. For instance, individuals with higher AQ scores show greater compensation for coarticulation in speech perception (Yu, 2010). In addition, higher AQ individuals are less sensitive to phonetic duration when assigning judgments of prosodic prominence (Bishop, 2016). Both sets of results suggest that higher AQ individuals rely more on system-level expectations (e.g., expected cooccurrence of phonetic signals), and less on the real-time signal itself (e.g., the phonetic duration of a perceived word).
@@
In order to apply this hypothesis to the case of contextual influence on the interpretation of have, we varied from 0.3 to 0.7 in steps of 0.1, and at each level, we ran 1000 simulations in each condition. As seen in Figure 10, higher values of corresponded with a reduction in the difference in acceptability between conditions. In other words, stronger field coupling reduced the magnitude of the contextual modulation effect. This is because input from the low CA distribution equally favors the low conn and high conn distributions. When this input is stronger, the lingering asymmetry between these distributions from phase 1 is reduced more quickly, reducing the influence of context on interpretation in phase 3.
By modeling the timecourse of individual instances of semantic interpretation, we can examine the relationship between interpretation content (the location of field stabilization) and interpretation time (the timestep of field stabilization) for individual simulations. As seen in Figure 11, this relationship was modulated by context. In the adjacency context, a negative correlation was observed: faster response times corresponded with higher acceptability ratings. This is because the adjacency context primed a more acceptable interpretation: when the field stabilized at a lower conn value (higher acceptability), this stabilization was facilitated by the context. In the possession context, a positive correlation was observed: faster response times corresponded with lower acceptability ratings. This is because the interpretation primed by the possession context was ultimately a less acceptable one.
@@
To date, studies of contextual facilitation of have interpretation have not simultaneously collected data regarding acceptability and processing time. Below we report an experiment combining acceptability judgments and self-paced reading. The purpose of the experiment is twofold: (a) to replicate previous results regarding contextual facilitation of adjacency readings of have, as well as individual variation in the magnitude of this effect indexed by the AQ, and (b) to test the new model prediction represented in Figure 11.
3 Experiment design
3.1 Participants
@@
3.2 Materials
Each experimental stimulus consisted of a pair of sentences (a context sentence followed by a target sentence) conjoined by and. Every target sentence was designed to convey an adjacency interpretation of have, as in 3.2.
@@ 3.2 Materials
It is difficult to construe a possession interpretation of 3.2 because oak trees do not typically possess skateboards (whether alienably or inalienably). Every target sentence had the form “the [noun1] has a [noun2] that is [adj]”. [noun2] in the target sentence is the critical word at which an interpretation of have can be construed, since after reading [noun2], the participant has read have and both of its arguments. The relative clause “that is [adj]” was included as a spillover region. There were ten target sentences, each of which was preceded by two different context sentences, for a total of 20 experimental stimuli. Each context sentence conveyed either an adjacency reading or an inalienable possession reading of have, as in Table 3.
@@
Every context sentence had the structure “the [noun1] has a [noun2] [modifier]”. [modifier] was either a prepositional phrase (in the adjacency condition) or a “that is [adj]” phrase (in the possession condition). [noun1] was identical between the two conditions in each set, and always contrasted saliently with [noun1] in the target sentence in order to increase overall felicitousness. In both conditions, [noun2] in the context was semantically unrelated to [noun2] in the target sentence. Moreover, the first phoneme in [noun2] in the context was always different from the first phoneme of [noun2] in the target sentence, in order to minimize confounds from phonological priming. All nouns were inanimate in order to maximize the availability of adjacency interpretations. 60 filler stimuli were also included. 20 of the fillers were of the same form as the experimental stimuli, but with contexts that used verbs other than have to convey an adjacency reading (10 stimuli) or a possession reading (10 stimuli). 40 of the fillers were completely unrelated to the experimental stimuli: 20 conveyed interpretations of circumstantial metonymy (e.g., “the grilled cheese at Table 6 ordered another coffee:), and 20 conveyed non-metonymous counterparts (e.g., “the customer at Table 6 ordered another coffee”). This yielded a total of 80 stimuli (20 experimental + 60 fillers).
@@
3.3 Procedure
Participants were instructed to complete the experiment in a quiet room where they would be free from distractions for up to an hour. To begin each trial, participants clicked on a button at the top of the screen with the words “Click here to begin the next trial”. Then, the first word of the stimulus appeared in the center of the screen. Participants pressed the spacebar to advance to the next word. Participants were instructed to read as quickly as possible while making sure to comprehend what they were reading. After participants advanced past the last word, they were prompted by the following instruction to give an acceptability rating of the entire stimulus: “How likely would you be to say this sentence, or hear this sentence from another native speaker of English?”. Participants gave their response on a seven-point Likert scale (labeled “very unnatural” on the left end and “very natural” on the right) by clicking on the corresponding button at the bottom of the screen.
The experiment began with four practice trials unrelated to the experimental stimuli. Then, each stimulus was presented twice, for a total of 160 trials per subject. Each of the two blocks was pseudo-randomized such that no two consecutive trials were from the same condition or the same stimulus set. The presentation order of the two blocks was counter-balanced between participants. After completing the combined self-paced reading and acceptability judgment task, participants completed the AQ. The entire procedure was conducted in the same session in Gorilla (Anwyl-Irvine et al., 2020). The session lasted approximately 30-40 minutes.
3.4 Data processing
@@
Reading times were log-transformed in order to approximate a normal distribution. In order to assess the effects of control variables on reading time, a linear mixed effects model was fit to the log-transformed RT (logRT) of all words with fixed effects of word length (in characters), trial number, and the preceding word’s logRT (all scaled and centered), a random intercept by participant, and random slopes for all predictors by participant. The model results are displayed in Table 4. All three control variables were found to significantly affect logRT: words took longer to read when they had more characters or when the preceding word took longer to read. Words were read faster when the word came later in the experiment. Thus, rather than analyzing raw logRT, we analyze the residuals of logRT after being regressed, by participant, on the three control predictors described above. In other words, we analyze the variance not predicted by the three control predictors.
@@
4 Experiment results
4.1 Acceptability
@@
Figure 12 displays mean z-scored (by-participant) acceptability ratings in each context. Both means are below 0, suggesting that the experimental stimuli were generally less acceptable than the filler stimuli, consistent with the general markedness of adjacency interpretations of have. However, consistent with the results of the regression model, ratings were higher in the adjacency context relative to the possession context.
Recall from Section 2 that in the model, context affected interpretation not by shifting the overall distribution of acceptability, but by increasing the likelihood of a distinct, higher-acceptability interpretation. We compare this simulated behavior to the measured behavior by examining the distribution of ratings in each context. Figure 13 demonstrates that, consistent with the simulations, the increase in mean acceptability in the adjacency context relative to the possession context was not accomplished via on overall shift in the distribution of ratings. For instance, the number of intermediate ratings 3 and 4 was largely unaffected by context. Rather, the increase in mean acceptability occurred because of a decrease in low ratings 1 and 2, and an increase in high ratings 5 and 6.
@@
In order to assess whether individual variation in the magnitude of contextual facilitation is predicted by AQ scores, we plot the by-participant random slopes for the effect of context against AQ score in Figure 14. A Spearman test confirms a negative correlation (, ): subjects with higher AQ scores showed a smaller effect of context, i.e. reduced contextual facilitation of adjacency have sentences.
4.2 Relationship between acceptability and reading time
@@
Next, we turn to the by-trial relationship between acceptability ratings and reading times. Recall that the model simulations predicted this relationship to be modified by context: a negative correlation in the adjacency context, and a positive correlation in the possession context. To test this prediction, we fit nested linear mixed effects models to the summed residualized reading times of the critical word [noun2] and each of the words in the spillover region “that is [adj]”. All models included random intercepts by item and by participant. Random slopes were not included because they led to model convergence issues and were also not of theoretical interest in this case.
Adding a fixed factor of context did not lead to a significant improvement in fit over a baseline model with only random effects (, ). However, adding a fixed factor for acceptability rating (z-scored by subject) did lead to a significant improvement in fit over the baseline model (, ). Adding a fixed factor for context back in did not lead to an improvement in fit over the model with only a fixed factor for acceptability (, ). Importantly, however, adding an interaction term did lead to a significant improvement in fit over the model with only fixed factors for main effects (, ). We report the results of the full model with both main effects and their interaction in Table 6. We observed a significant negative interaction between context and acceptability, such that the relationship between acceptability and reading time was more negative in the adjacency context relative to the possession context. This result, also observable in Figure 15, is consistent with the model prediction described in Section 2.2.
@@ 4.3 Discussion
The acceptability results confirm previously reported results regarding contextual facilitation of adjacency have interpretations. The acceptability of adjacency have sentences was improved following a context sentence that also conveyed an adjacency have interpretation, relative to a context sentence that conveyed a possession have interpretation. Moreover, individual variation in the magnitude of this improvement was predicted by AQ scores: individuals with higher AQ scores (lower context-sensitivity) showed reduced contextual facilitation. Finally, consistent with the novel prediction of the model simulations, the by-trial relationship between reading time and acceptability evidenced context modulation. The relationship was more negative in the adjacency context relative to the possession context.

### v2

@@ 2.4 Simulated individual variation
Previous empirical results have also demonstrated that individual variation in the magnitude of the by-context difference in acceptability is predicted by AQ scores, such that individuals with higher AQ scores show a reduced influence of context (Zhang et al., 2022). We model individual variation in AQ by varying the parameter , which controls the magnitude of field-field coupling. Stronger field-field coupling is consistent with a greater degree of system-level expectations. In other words, given some interpretation on one semantic dimension (e.g., control asymmetry (CA)), individuals can vary in the degree to which they expect a corresponding interpretation on a related semantic dimension (e.g., connectedness (conn)). We posit that individuals with higher AQ scores are more influenced by system-level expectations, i.e. stronger field-field coupling. With stronger system-level expectations, the processing system is more rigid, and less influenced by real-time information. This is consistent with existing findings relating AQ scores to linguistic behavior. For example, individuals with higher AQ scores show greater compensation for coarticulation in speech perception (Yu, 2010). In addition, higher AQ individuals are less sensitive to phonetic duration when assigning judgments of prosodic prominence (Bishop, 2016). Both sets of results suggest that higher AQ individuals rely more on system-level expectations (e.g., expected cooccurrence of phonetic signals), and less on the real-time signal itself (e.g., the phonetic duration of a perceived word). Stronger excitatory coupling between neural populations is also consistent with the hypothesis that autistic traits correlate with an increased ratio of neural excitation/inhibition (Rubenstein and Merzenich, 2003).
@@
In order to apply this hypothesis to the case of contextual influence on the interpretation of have, we varied from 0.25 to 0.45 in steps of 0.05, and at each level, we ran 1000 simulations in each condition. As seen in Figure 12, higher values of corresponded with a reduction in the difference in acceptability between conditions. In other words, stronger field-field coupling reduced the magnitude of the contextual modulation effect. This is because input from the low CA distribution equally favors the low conn and high conn distributions. When this input is stronger, the lingering asymmetry between these distributions from phase 1 is reduced more quickly, thus reducing the influence of context on interpretation in phase 3.
2.5 Simulated response time
Next, we examine the influence of context on response time (RT). We operationalize RT in each simulation as the number of timesteps between the onset of phase 3 and the timestep at which the first neuron passed the threshold for lateral interaction (), leading to activation peak stabilization. Figure 13 displays mean RT by context.
Consistent with self-paced reading data (Zhang et al., 2018), RT was slower in the possession context relative to the adjacency context. In the model, the reason for this is as follows. Across contexts, when an adjacency interpretation occurs (activation peak at low conn), this tends to occur faster (lower RT) than when a possession interpretation occurs (high conn), because of the influence of the weak input to low conn from the nouns. At the same time, the activation peak is more likely to form at low conn in the adjacency context relative to the possession context, due to the biasing influence of phase 1 activation on phase 3 activation. Taken together, this leads to overall slower RT in the possession context relative to the adjacency context.
This combination of biases has another interesting consequence: an interaction between context and acceptability in predicting RT, as seen in Figure 14.
As mentioned above, across contexts, low conn (higher acceptability) interpretations tend to be reached more quickly than high conn (lower acceptability) interpretations, due to the bias from the nouns (weak low conn input). This leads to an overall negative correlation between acceptability and RT. Moreover, in the adjacency context, activation from phase 1 (the context sentence) biases activation in phase 3 (the target sentence) towards low conn (higher acceptability), so that low conn peaks form even more quickly. This reinforces the overall negative correlation between acceptability and RT in the adjacency context, as seen in Figure 14 (). In the possession context, on the other hand, activation from phase 1 biases activation in phase 3 towards high conn (lower acceptability), competing with the overall bias towards low conn from the nouns. The outcome of this competition is that the correlation between acceptability and RT in the possession context is reversed (). Notably, the positive correlation in the possession context is of a weaker magnitude than the negative correlation in the adjacency context, since the correlation in the adjacency context is the result of two reinforcing influences, while the correlation in the possession context is the outcome of two competing influences.
This account is supported by a linear regression model of simulated RT by simulated acceptability (z-scored) and context (sum-coded: possession , adjacency ). The model revealed main effects of context (, , ) and acceptability (, , ), and an interaction between context and acceptability (, , ). In order to confirm that the magnitude of the negative correlation in the adjacency context was greater than the magnitude of the positive correlation in the possession context, we multipled RT values in the possession context by , and then ran the same linear model as above. The interaction between context and acceptability remained significant (, , ), confirming a difference in the magnitudes of the correlations, independent of the directions.
3 Behavioral experiment
@@
To date, studies of contextual facilitation of have interpretation have not simultaneously collected data regarding acceptability and processing time. In this section, we report an experiment combining acceptability judgments and self-paced reading. The purpose of the experiment is twofold: (a) to replicate previous results regarding contextual facilitation of adjacency readings of have, as well as individual variation in the magnitude of this effect indexed by the AQ, and (b) to test the model prediction represented in Figure 14, which has not previously been tested.
3.1 Experiment design
3.1.1 Participants
@@
3.1.2 Materials
Each experimental stimulus consisted of a pair of sentences (a context sentence followed by a target sentence) conjoined by and.888All stimuli are included in A. Every target sentence was designed to convey an adjacency interpretation of have, as in 3.1.2.
@@ 3.1.2 Materials
It is difficult to construe a possession interpretation of 3.1.2 because oak trees do not typically possess skateboards (whether alienably or inalienably). Every target sentence had the form “the [noun1] has a [noun2] that is [adj]”. [noun2] in the target sentence is the critical word at which an interpretation of have can be construed, since after reading [noun2], the participant has read have and both of its arguments. The relative clause “that is [adj]” was included as a spillover region. There were ten target sentences, each of which was preceded by two different context sentences, for a total of 20 experimental stimuli. Each context sentence conveyed either an adjacency reading or an inalienable possession reading of have, as in Table 3.
@@
Every context sentence had the structure “the [noun1] has a [noun2] [modifier]”. [modifier] was either a prepositional phrase (in the adjacency condition) or a “that is [adj]” phrase (in the possession condition). [noun1] was identical between the two conditions in each set, and always contrasted saliently with [noun1] in the target sentence in order to increase overall felicitousness. In both conditions, [noun2] in the context was semantically unrelated to [noun2] in the target sentence. Moreover, the first phoneme in [noun2] in the context was always different from the first phoneme of [noun2] in the target sentence, in order to minimize confounds from phonological priming. All nouns were inanimate in order to maximize the availability of adjacency readings. 60 filler stimuli were also included. 20 of the fillers were of the same form as the experimental stimuli, but with contexts that used verbs other than have to convey an adjacency reading (10 stimuli) or a possession reading (10 stimuli). 40 of the fillers were completely unrelated to the experimental stimuli: 20 conveyed interpretations of circumstantial metonymy (e.g., “the grilled cheese at Table 6 ordered another coffee”), and 20 conveyed non-metonymous counterparts (e.g., “the customer at Table 6 ordered another coffee”). This yielded a total of 80 stimuli (20 experimental + 60 fillers).
@@
3.1.3 Procedure
Participants were instructed to complete the experiment in a quiet room where they would be free from distractions for up to an hour. To begin each trial, participants clicked on a button at the top of the screen with the words “Click here to begin the next trial”. Then, the first word of the stimulus appeared in the center of the screen. Participants pressed the spacebar to advance to the next word, i.e., word-by-word self-paced reading (Just et al., 1982). Participants were instructed to read as quickly as possible while making sure to comprehend what they were reading. After participants advanced past the last word, they were prompted by the following instruction to give an acceptability rating of the entire stimulus: “How likely would you be to say this sentence, or hear this sentence from another native speaker of English?”. Participants gave their response on a seven-point Likert scale (labeled “very unnatural” on the left end and “very natural” on the right) by clicking on the corresponding button at the bottom of the screen.
The experiment began with four practice trials unrelated to the experimental stimuli. Then, each stimulus was presented twice in order to increase robustness to noise, given that the data was collected online.999An anonymous reviewer raises the potential concern of unintended priming at the second presentation of each stimulus. In order to investigate this, we also conducted all analyses separately for each presentation block. The pattern of results reported in Section 3.2 was mostly equivalent between blocks. In particular, in both blocks, we observed a participant-level correlation between AQ score and context effect on acceptability (see Section 3.2.1), and an interaction between context and acceptability in predicting reading time (see Section 3.2.2). However, there was a difference between blocks in the relationship between context, trial number, and acceptability. The effect of context on acceptability was only significant in block 2 and not block 1, and the effect of trial number on acceptability was only significant in block 1 and not block 2. In order to investigate this pattern further, we ran a single linear mixed effects model with the same structure as the model described in Section 3.2.1, except that it also included a main effect of block, as well as all interactions between block, context, and trial number. Consistent with the model reported in Table 5, we found main effects of trial number (, , ) and context=adjacency (, , ). We also found a two-way interaction between context=adjacency and trial number (, , ), a two-way interaction between block=2 and trial number (, , ), and a marginal three-way interaction between context=adjacency, block=2, and trial number (, , ). This pattern of results suggests that acceptability ratings in the possession context generally decreased over the course of block 1, before reaching a minimum in block 2. Priming would not explain this effect, since priming should lead to an increase in acceptability in block 2 relative to block 1. Therefore, we interpret the difference between blocks as primarily the result of nonlinearity in the effect of trial number. In the remainder of the paper, we report results from across the two presentation blocks. Each of the two blocks was pseudo-randomized such that no two consecutive trials were from the same condition or the same stimulus set. The presentation order of the two blocks was counter-balanced between participants. Each participant completed 160 trials in total. After completing the combined self-paced reading and acceptability judgment task, participants completed the AQ. The entire procedure was conducted in the same session in Gorilla (Anwyl-Irvine et al., 2020). The session lasted approximately 30-40 minutes.
3.1.4 Data processing
@@
Word reading times were log-transformed in order to approximate a normal distribution. In order to assess the effects of control variables on reading time, a linear mixed effects model was fit to the log-transformed RT (logRT) of all words with fixed effects of word length (in characters), trial number, and the preceding word’s logRT (all scaled and centered), a random intercept by participant, and random slopes for all predictors by participant. The model results are displayed in Table 4. All three control variables were found to significantly affect logRT: words took longer to read when they had more characters or when the preceding word took longer to read. Words were read faster when the word came later in the experiment. Thus, rather than analyzing raw logRT, we analyze the residuals of logRT after being regressed, by participant, on the three control predictors described above. In other words, we analyze the variance not predicted by the three control predictors.
@@
3.2 Experiment results
3.2.1 Acceptability
@@
Figure 15 displays mean z-scored (by-participant) acceptability ratings in each context. Both means are below 0, suggesting that the experimental stimuli were generally less acceptable than the filler stimuli, consistent with the general markedness of the target sentences (Zhang et al., 2018), which was also evident in simulated acceptability (Section 2.3). Consistent with the results of the regression model, ratings were higher in the adjacency context relative to the possession context.
@@
In order to assess whether individual variation in the magnitude of contextual facilitation is predicted by AQ scores, we plot the by-participant slopes for the effect of context against AQ score in Figure 16. A Spearman test confirms a negative correlation (, ): subjects with higher AQ scores showed a smaller effect of context, i.e. reduced contextual facilitation of adjacency have sentences.
This trend can also be seen in Figure 17, which plots mean rating (z-scored by participant) by condition against participant AQ scores, binned into quintiles, analogous to Figure 12 which displays the simulation results. While not as clean as the simulation results, there is a general trend that the difference in mean rating between conditions tends to decrease as AQ score increases.
3.2.2 Reading time
@@
Next, we turn to the reading time results. We fit nested linear mixed effects models to the summed residualized log-transformed reading times (see Section 3.1.4) of the critical word [noun2] and each of the words in the spillover region “that is [adj]”. All models included random intercepts by item and by participant.101010Random slopes were not included because they led to model convergence issues and were also not of theoretical interest in this case. Using the same methods for fitting nested linear mixed effects models as above (Section 3.2.1), we assessed the influence of the following fixed factors on reading time: context (sum-coded: possession , adjacency ), acceptability (z-scored by subject), and their interaction. Adding a fixed factor of context did not lead to a significant improvement in fit over a baseline model with only random effects (, ). However, adding a fixed factor for acceptability rating did lead to a significant improvement in fit over the baseline model (, ). Adding a fixed factor for context back in did not lead to an improvement in fit over the model with only a fixed factor for acceptability (, ). Importantly, however, adding an interaction term did lead to a significant improvement in fit over the model with only fixed factors for main effects (, ). We report the results of the full model with both main effects and their interaction in Table 6.
@@
Consistent with the model simulation results described in Section 2.5, we observed a main negative effect of acceptability on reading time. Also consistent with the simulation results, we observed an interaction between context and acceptability in predicting reading time: the relationship between acceptability and reading time was more negative in the adjacency context relative to the possession context (Figure 18).
Finally, as predicted by the model, we observed a significant negative correlation between acceptability and reading time in the adjacency context ().
Two effects predicted by the model were not observed. First, we did not observe the predicted positive correlation between acceptability and reading time in the possession context (). We discuss a possible explanation of this discrepancy between prediction and observation in Section 4.1. Second, we did not observe a main effect of context on reading time (Figure 19).
This result contrasts with the prediction from the simulations, as well as previous results (Zhang et al., 2018). In order to investigate this discrepancy, we examined the reading time of each word in the critical region, displayed in Figure 20.
We fit separate linear mixed effects models to residualized log-transformed reading time for each word, with the same structure as the model reported in Table 6. There was no main effect of context on RT for the critical word [noun2] (, , ) or the following word “that” (, , ). However, at the verb “is” (two words after the critical word), there was a main effect of context on RT in the opposite direction to the prediction (, , ). Not until [adj], three words after the critical word, did we observe a main effect of context on RT in the predicted direction (, , ). We hypothesize that the unpredicted effect on “is” was due to unintended lexical priming in the possession context, since both the possession context sentence and the target sentence included a relative clause beginning with “that is”, while the adjacency context sentence did not include this phrase. This low-level similarity between the possession context sentence and the target sentence, which was not present in the stimuli of Zhang et al. (2018), likely obscured the main effect of context on RT when measured across the entire critical region. This said, our replication of the acceptability effect and the by-participant correlation between the acceptability effect and AQ score, as well as our confirmation of the predicted interaction between context and acceptability in predicting reading time, suggests that this difference in stimulus design did not otherwise confound the results of the experiment.
@@ 3.3 Discussion
The acceptability results replicate previously reported results regarding contextual facilitation of adjacency have readings. Specifically, the acceptability of sentences with have and two unrelated, inanimate nouns was improved following a context sentence that conveyed an adjacency have reading, relative to a context sentence that conveyed a possession have reading (Figure 15). Moreover, individual variation in the magnitude of this improvement was predicted by AQ scores: individuals with higher AQ scores (lower context-sensitivity) showed reduced contextual facilitation (Figures 16, 17).
With respect to reading time, the predicted main effect of context on reading time was observed on the last word in the critical region; at the previous word, the opposite effect was observed (Figure 20), arguably due to unintended priming in the possession context, since both the possession context sentence and the target sentence contained the phrase “that is”. Finally, the predicted relationship between context, acceptability, and reading time was partially supported (Table 6, Figure 18). As predicted, there was a significant main effect of acceptability on reading time, a significant interaction between context and acceptability in predicting reading time, and a significant negative correlation between reading time and acceptability in the adjacency context. However, the predicted positive correlation between reading time and acceptability in the possession context was not observed. We discuss a possible explanation for this unpredicted result in the following section.